OpenStudy (anonymous):

The elephant population in American zoos is declining at a .15% rate. If the elephant population was 2312 in 1995, use the exponential equation to predict the population in 2015.

OpenStudy (rulnick):

It is a .15% rate and not a 15% rate, right? Then the prediction would be 2312(1-0.0015)^(2015-1995) = 2312(0.9985)^20 = ...

OpenStudy (rulnick):

Need more help?

OpenStudy (rulnick):

If so: (0.9985)^20 = 0.970424 (rounded). So just multiply this by 2312 for your answer.